OpenStudy (anonymous):

Solve for x. Write answer to three decimal places. (1000)(1.05)^3x = 3000. Will award medal!!!!

OpenStudy (anonymous):

x=2.5915 soved for X now you have move decimal not sure?????

OpenStudy (♪chibiterasu):

@binks Please don't answer questions if you are not going to guide the asker through it, that doesn't help them learn. Now let's see: We have (1000(1.05) which means we multiply them, so what do we get when we multiply 1000 by 1.05?

OpenStudy (♪chibiterasu):

Do do do do, do do do do, do do do do . . . DING! The product of said numbers is 1050 Now we have 1050^3x = 3000

OpenStudy (anonymous):

ok

OpenStudy (♪chibiterasu):

Hm, I just noticed that 3x is raised . . .

OpenStudy (anonymous):

yes

OpenStudy (♪chibiterasu):

Well I'm not so sure about the 3x but I assume you meant this: \[(1000)(1.05)^{3x} = 3000\]

OpenStudy (anonymous):

yes that is what i meant

OpenStudy (anonymous):

sorry I just wasn't sure aout her question. that's why I asked

OpenStudy (♪chibiterasu):

Well please don't give straight answers, this is Open "Study". We'd rather you guide people through their questions.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

can someone explain how to get the answer and show the final answer?

OpenStudy (♪chibiterasu):

@binks apparently knows the answer, so he should be able to help you :3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

actually 2.5915 is incorrect. You will log both sides. 3xLog 1.05=Log3. Then you will divide both sides by Log 1.05 ... 3x(Log1.05/Log1.05)=Log3/Log1.05. this will simplify to 3x=Log3/Log1.05. Then use your calculator to simplify down to x=7.506